Please put it excel spreadsheet format and please provide answers and formula. Thank you.

LU 8-2 PRACTICE QUIZ ive the followding situations (markups based on selling price). Note numbers 1,2 are parallel problems to those in Practice Quiz 8-1 lowing situatio omplete this Practice Qutz to ee how you are dotng 1. Irene Westing bought a desk for $400 from an office supply house. She plans to sellthe desk for $600. What is Irenes dollar markup? What is her percent markup on selino price rounded to the nearest tenth percent)? Check your answer. Selling price will e slightly off due to rounding Suki Komar bought dolls for her toy store that cost $12 each. To make her desired profit, Sukimust mark up each doll 35% on the selling price. What is the selling price of each doll? What is the dollar markup? Check your answer Jay Lyman sells calculators. His competitor sells a new calculator line for $14 each. Ja needs a 40% markup on the selling price to make his desired profit, and he must meet price competition. What is Jays dollar markup? At what cost can Jay afford to bring these calculators into the store? Check your answer Dan Flow sells wrenches for $10 that cost $6. What is Dans percent markup on cost? Round to the nearest tenth percent. What is Dans percent markup on selling price? Check your answer 2. 3. 4. LU 8-2a TRA PRACTICE QUIZ WITH WORK Solve the following situations (markups based on selling price). 1. Irene Westing bought a desk for $800 from an office supply house. She plans to this Need more practice? Try Extra Practice Qulz (check sell the ing price willbe her desite Chapter desk for $1,200. What is Irenes dollar markup? What is her percent markup on price (rounded to the nearest tenth percent)? Check your answer. Selling price selily slightly off due to rounding t Solutions be found tn Appendix B. 2. Suki Komar bought dolls for her toy store that cost $14 each. To make ice. What is the selling price of profit, Suki m ust mark up each doll 38% on selling pr each doll? What is the dollar markup? Check your answer. Jay Lyman sells calculators. His competitor sells a new calculator line for $16 each,. la needs a 42% markup on the selling price to make his desired profit, and he must meet price competition. What is Jays dollar markup? At what cost can Jay afford to bring these calculators into the store? Check your answer Dan Flow sells wrenches for $12 that cost $7. What is Dans percent markup on cost? Round to the nearest tenth percent. What is Dans percent markup on selling price? Check your answer 3. 4.
Answer #1

As per rules I am answering the first 4 subparts of the question

1: Dollar markup = Selling price - cost = 600-400 = $200

Percentage markup = Mark up/ SP

= 200/600 = 33.33%

2

Cost 12
Mark upon SP 35%
Mark up on cost 53.85%
Mark up in dollar 6.46
Selling price 18.46

3:

SP 14
Mark up 40%
Dollar mark up 5.6
Cost 8.4

4:

Cost 6
SP 10
Mark up 4
%Mark up on cost 66.67%
% Mark up on sp 40%

WORKINGS

AutoSave Book1 Excel (Product Activation Failed) Sign in File Home Insert Draw Page Layout Formulas Data Review View Help Tell me what you want to do 수 Share Calibri Custom Conditional Fornat as Cell Insert Delete Fornat FormattingTable Styles- 、. Sort & Find & Paste . 녀 . 의 _ . 로三들經垣臣 Merge & Center-5 . % , : Filter-Select- Clipboard Alignment Number Styles Cells Editing B2 ﹀丘1/1/1900 12:00:00 AM Cost SP Mark up 400 600 D3-D2 D4/D3 3 SP 14 0.4 Mark up Dollar mark up 40 14/100 Cost -H2-H4 4 Cost Cost Mark upon SP 0.35 Mark up on cost :35%*100/65 Mark up in dollar D9 D7 Selling price 12 SP Mark up %Mark up on cost -H3/H7 % Mark up on sp -H3/HB 10 H8-H7 10 D7+D10 12 13 15 16 17 Sheet1 Ready + 100